# Staying on Top of SMEFT-Likelihood Analyses

### v1

@@ 5 Outlook
Global SMEFT analyses are an exciting development at the LHC, as they combine its role as a precision hadron collider with the goal of interpreting all measurements in terms of precision quantum field theory. This precision theme implies that even if we know that the current measurements do not rule out the Standard Model, limits on SMEFT Wilson coefficients tell us important information about fundamental physics.
To extract limits on fundamental physics parameters we need a comprehensive uncertainty treatment, covering experimental statistical uncertainties, experimental systematics, and theory uncertainties. For the latter two, it is crucial that we include correlations. Public likelihoods are the state of the art in communicating such experimental results to a broader community. We include, for the first time, public ATLAS likelihoods for cross section measurements in a global analysis. These public likelihoods allow us to systematically evaluate the effects of correlations of systematic and theory uncertainties on a global analysis.
The basis of the global SFitter analysis is a fully exclusive likelihood. It includes a large set of rate and kinematic measurements, either pre-processed by ATLAS or CMS, unfolded, or extracted and backward-engineered from experimental publications. The uncertainty treatment is especially flexible, including a choice of flat nuisance parameters for correlated theory uncertainties. Starting from the fully exclusive likelihood we can employ a profile likelihood or a Bayesian marginalization to extract limits on individual Wilson coefficients. In the top sector, we find no significant difference between the two statistical approaches.
A focus of this paper was on the role of different correlated uncertainties. In a similar analysis, albeit without public likelihoods, we found that in the electroweak sector the correlations were crucial, whereas the theory uncertainties were not (yet) a limiting factor [65]. Intriguingly, the situation in the top sector is the opposite: theory uncertainties are crucial, while the correlations of experimental systematics have limited impact on the SMEFT limits. This reflects the QCD nature and the vast statistics of top pair production.
We emphasize that this is not a final statement about public likelihoods in SMEFT analyses. The reason is that we find kinematic measurements of boosted top pair production to be the driver behind improved SMEFT limits. For unfolded kinematic measurements there are no public likelihoods available, but we are looking forward to implementing them in SFitter in the future.
@@ Acknowledgements
We would like to thank Dirk Zerwas, James Moore and Luca Mantani for discussions on the challenges of Monte Carlo toys, as well as Sabine Kraml and Lukas Heinrich for the encouragement to use public likelihoods. We are also grateful to Luca Mantani for his help with the calculation of SMEFT predictions and Tomas Dado for his hands-one help with public likelihoods. This research is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257 Particle Physics Phenomenology after the Higgs Discovery. NE is funded by the Heidelberg IMPRS Precision Tests of Fundamental Symmetries. We acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 39/963-1 FUGG (bwForCluster NEMO).
@@ Appendix A Toys vs Markov chain
The current SFitter methodology relies on a Markov chain to encode the fully exclusive likelihood given in Eq.(8). It then allows for a profile likelihood or a marginalization to remove nuisance parameters and extract limits on individual Wilson coefficients [63, 65]. Past SFitter analyses used an alternative methodology known as Monte Carlo toys [63, 64, 12] or Monte Carlo replicas [134, 21, 138, 39]. This method has not been used in SFitter since the SMEFT analysis of the top sector [22]. In Ref. [39], its shortcomings for a SMEFT analysis of the top sector are discussed in detail. In this appendix we provide an additional discussion of Monte Carlo toys in the SFitter context.
@@
The toys method describes possible outcomes of a measurement, given its uncertainty, with a nuisance parameter. In the Gaussian limit the outcome of a measurement is sampled around the mean $\bar{d}$,
@@
An interesting challenge is the description of correlations between measurements in the likelihood. For the Gaussian case, we can describe them with a correlation matrix, if the correlation is less than 100%. To describe correlations for non-Gaussian likelihoods, we introduce nuisance parameters, which increase the effective dimensionality of the likelihood.
@@
Figure 14: Comparison between toys and MCMC, constraining $c_{1}$ and $c_{2}$ using the measurement $d_{1}$ from Eq.(16). The 1-dimensional distributions show the profile likelihoods as well as the marginalized probabilities. The profiling is obtained by analytically maximizing the log-likelihood of Eq.(12) along each axis.
@@
This relation cannot be solved for real $c_{i}$, and the spread from $\delta_{1,2}=0.3$ is too small to cure this problem. The likelihoods obtained from toys and from the Markov chain are both given in the left panel of Fig. 15, and they differ. The Markov chain includes points with a finite likelihood offset below the theoretical maximum. The toys are derived from Eq.(14), returning $c_{1}=c_{2}=0$ if they cannot reach the true maximum. For the right panel of Fig. 15 we modify the toys algorithm to remove the maximum-likelihood peak at the parameter space boundary. This is done by retaining only the samples $d_{k}$ in Eq.(13) which satisfy $d_{k}>m_{\text{SM,2}}$. With this modification the Markov chain and the toys give the same likelihood.
@@
Summarizing Fig. 15, the toys and the Markov chain treat the unwanted parameter space $c_{i}^{2}<0$ differently. While the toys provide a maximum-likelihood parameter point for each assumed measurement, the Markov chain removes the unwanted points. In terms of a prior on the numerical implementation of the scanning we can understand the two methods as
@@
where $\Delta$ is chosen to remove the numerically broadened peak at $c=0$. In the right panel of Fig. 15 we see that for allowed $c_{i}$ the two methods give the same result, but with a different normalization of the probability $p(c|d)$ [39]. If we work with the likelihood $p(d|c)$, the normalization does not matter. If we include all toy experiments, the two distributions in Fig. 15 reflect the fact that the two methods are asking different statistical questions.
@@
The situation given by Eq.(17) is just an extremely unlucky outcome. Of very few measurements one happens to be many standard deviations away from the prediction, while we expect such an outlier only once in many more measurements.
Moreover, in a realistic global analysis, many measurements together constrain a given model parameter. To see what happens then, we first combine $d_{1}$ and $d_{2}$ in the left panel of Fig. 16. The Markov chain generates a valid point distribution, symmetric in ±$\pm|c_{2}|$. For the toys, the situation is different. On the underfluctuation in $d_{2}$, $c_{1}$ and $c_{2}$ have the same impact, but for the overfluctuation in $d_{1}$ a shift in $c_{2}$ gains more. This is why a peak is observed at $c_{1}=0$ and $c_{2}$ is adjusted.
@@
The central maximum of the corresponding likelihood will be at $c_{1}\approx 1.6$. In the right panel of Fig. 16 we see that the peak at $c_{1}=0$ is now accompanied by a distribution of finite $c_{1}$ symmetric around zero. This way, the importance of the peak is reduced. For more measurements this will continue, until the underfluctuation will just be an expected statistical outlier in a large set of measurements, with little effect on the global likelihood.
@@
Because our measurements show a purely quadratic dependence, we can try to circumvent our problem by extracting the likelihood over a model space defined by $c_{1}^{2}$ and $c_{2}^{2}$. Instead of first requiring $c_{i}^{2}$ to be positive and then maximizing the log-likelihood, we first maximize the log-likelihood in terms of the $c_{i}^{2}$. In the left panel of Fig. 17 we show the results in terms of the $c_{i}^{2}$, with toys and Markov chain in perfect agreement. However, almost the entire preferred range gives unwanted values for $c_{2}^{2}$. The curve defined by $\Delta\chi^{2}=16$ highlights how unlikely it is to obtain points with $c_{i}^{2}>0$.
In a second step, we now select valid parameter points. This avoids a hard boundary when optimizing the toys. In the right panel of Fig. 17, the blue curves show the points in Markov chain, sampling directly in $c_{1}$ vs $c_{2}$. These are the same points as those in the upper-right quadrant of the left panel of Fig. 16. In orange we show the one-dimensional distributions of only the allowed parameter points, out of all those plotted in the left panel of Fig. 17. Neither the likelihood nor the probability is parametrization invariant, so we need to apply the Jacobian

### v3

@@ 5 Outlook
Global SMEFT analyses are an exciting development at the LHC, as they combine their role as a precision hadron collider with the goal of interpreting all measurements in terms of precision quantum field theory. This precision theme implies that even if we know that the current measurements do not rule out the Standard Model, limits on SMEFT Wilson coefficients tell us important information about fundamental physics.
To extract limits on fundamental physics parameters, we need a comprehensive uncertainty treatment covering experimental statistical uncertainties, experimental systematics, and theory uncertainties. For the latter two, it is crucial that we include correlations. Public likelihoods are the state of the art in communicating such experimental results to a broader community. We include, for the first time, public ATLAS likelihoods for cross section measurements in a global analysis. These public likelihoods allow us to systematically evaluate the effects of correlations of systematic and theory uncertainties on a global analysis.
The basis of the global SFitter analysis is a fully exclusive likelihood. It includes a large set of rate and kinematic measurements, either pre-processed by ATLAS or CMS, unfolded, or extracted and backward-engineered from experimental publications. The uncertainty treatment is especially flexible, including a choice of flat nuisance parameters for correlated theory uncertainties. Starting from the fully exclusive likelihood, we can employ a profile likelihood or a Bayesian marginalization to extract limits on individual Wilson coefficients. In the top sector, we find no significant difference between the two statistical approaches.
The focus of this paper was on the role of different uncertainties, their correlations and the role of public likelihoods in this context. In a similar analysis, albeit without public likelihoods, we found that in the electroweak sector, the correlations were crucial, whereas the theory uncertainties were not (yet) a limiting factor [65]. Intriguingly, the situation in the top sector is the opposite: theory uncertainties are crucial, while the correlations of experimental systematics have a limited impact on the SMEFT limits. This reflects the QCD nature and the vast statistics of top pair production.
We have demonstrated that public likelihoods provide a much more flexible approach to handling nuisance parameters. However, fully leveraging their potential currently proves difficult due to the large number of measurements included in our global analysis. We emphasize that this is not a final statement about public likelihoods in SMEFT analyses. The reason is that we find kinematic measurements of boosted top pair production to be the driver behind improved SMEFT limits. For unfolded kinematic measurements, there are no public likelihoods available yet, but we are looking forward to implementing them in SFitter in the future.
We finished this study of the impact of theory uncertainties in a consistent theory framework of LHC data by performing the first combined SFitter analysis of the Higgs, electroweak, and top sectors. This further displayed the limiting effect of theory uncertainties on the constraining power of modern top measurements compared to those in the Higgs sector.
@@ Acknowledgements
We would like to thank Dirk Zerwas, James Moore, and Luca Mantani for discussions on the challenges of Monte Carlo toys, as well as Sabine Kraml and Lukas Heinrich for the encouragement to use public likelihoods. We are also grateful to Luca Mantani for his help with the calculation of SMEFT predictions and Tomas Dado for his hands-on help with public likelihoods. This research is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257 Particle Physics Phenomenology after the Higgs Discovery. NE is funded by the Heidelberg IMPRS Precision Tests of Fundamental Symmetries. MM acknowledges support from the Alexander von Humboldt Foundation. We acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 39/963-1 FUGG (bwForCluster NEMO).
@@ Appendix A Toys vs Markov chain
The current SFitter methodology relies on a Markov chain to encode the fully exclusive likelihood given in Eq.(8). It then allows for a profile likelihood or marginalization to remove nuisance parameters and extract limits on individual Wilson coefficients [63, 65]. Past SFitter analyses used an alternative methodology known as Monte Carlo toys [63, 64, 12] or Monte Carlo replicas [134, 21, 147, 39]. This method has not been used in SFitter since the SMEFT analysis of the top sector [22]. In Ref. [39], its shortcomings for a SMEFT analysis of the top sector are discussed in detail. In this appendix, we provide an additional discussion of Monte Carlo toys in the SFitter context.
@@
The toys method describes possible outcomes of a measurement, given its uncertainty, with a nuisance parameter. In the Gaussian limit, the outcome of a measurement is sampled around the mean $\bar{d}$,
@@
An interesting challenge is the description of correlations between measurements in the likelihood. For the Gaussian case, we can describe them with a correlation matrix if the correlation is less than 100%. To describe correlations for non-Gaussian likelihoods, we introduce nuisance parameters, which increase the effective dimensionality of the likelihood.
@@
Figure 15: Comparison between toys and MCMC, constraining $c_{1}$ and $c_{2}$ using the measurement $d_{1}$ from Eq.(17). The 1-dimensional distributions show the profile likelihoods, as well as the marginalized probabilities. The profiling is obtained by analytically maximizing the log-likelihood of Eq.(13) along each axis.
@@
This relation cannot be solved for real $c_{i}$, and the spread from $\delta_{1,2}=0.3$ is too small to cure this problem. The likelihoods obtained from toys and from the Markov chain are both given in the left panel of Fig. 16, and they differ. The Markov chain includes points with a finite likelihood offset below the theoretical maximum. The toys are derived from Eq.(15), returning $c_{1}=c_{2}=0$ if they cannot reach the true maximum. For the right panel of Fig. 16, we modify the toys algorithm to remove the maximum-likelihood peak at the parameter space boundary. This is done by retaining only the samples $d_{k}$ in Eq.(14) which satisfy $d_{k}>m_{\text{SM,2}}$. With this modification, the Markov chain and the toys have the same likelihood.
@@
Summarizing Fig. 16, the toys and the Markov chain treat the unwanted parameter space $c_{i}^{2}<0$ differently. While the toys provide a maximum-likelihood parameter point for each assumed measurement, the Markov chain removes the unwanted points. In terms of a prior on the numerical implementation of the scanning, we can understand the two methods as
@@
where $\Delta$ is chosen to remove the numerically broadened peak at $c=0$. In the right panel of Fig. 16 we see that for allowed $c_{i}$ the two methods give the same result but with a different normalization of the probability $p(c|d)$ [39]. If we work with the likelihood $p(d|c)$, the normalization does not matter. If we include all toy experiments, the two distributions in Fig. 16 reflect the fact that the two methods are asking different statistical questions.
@@
The situation given by Eq.(18) is just an extremely unlucky outcome. Of very few measurements, one happens to be many standard deviations away from the prediction, while we expect such an outlier only once in many more measurements.
Moreover, in a realistic global analysis, many measurements together constrain a given model parameter. To see what happens then, we first combine $d_{1}$ and $d_{2}$ in the left panel of Fig. 17. The Markov chain generates a valid point distribution, symmetric in ±$\pm|c_{2}|$. For the toys, the situation is different. On the underfluctuation in $d_{2}$, $c_{1}$ and $c_{2}$ have the same impact, but for the overfluctuation in $d_{1}$, a shift in $c_{2}$ gains more. This is why a peak is observed at $c_{1}=0$ and $c_{2}$ is adjusted.
@@
The central maximum of the corresponding likelihood will be at $c_{1}\approx 1.6$. In the right panel of Fig. 17, we see that the peak at $c_{1}=0$ is now accompanied by a distribution of finite $c_{1}$ symmetric around zero. This way, the importance of the peak is reduced. For more measurements, this will continue until the underfluctuation will just be an expected statistical outlier in a large set of measurements, with little effect on the global likelihood.
@@
Because our measurements show a purely quadratic dependence, we can try to circumvent our problem by extracting the likelihood over a model space defined by $c_{1}^{2}$ and $c_{2}^{2}$. Instead of first requiring $c_{i}^{2}$ to be positive and then maximizing the log-likelihood, we first maximize the log-likelihood in terms of the $c_{i}^{2}$. In the left panel of Fig. 18, we show the results in terms of the $c_{i}^{2}$, with toys and Markov chain in perfect agreement. However, almost the entire preferred range gives unwanted values for $c_{2}^{2}$. The curve defined by $\Delta\chi^{2}=16$ highlights how unlikely it is to obtain points with $c_{i}^{2}>0$.
In the second step, we now select valid parameter points. This avoids a hard boundary when optimizing the toys. In the right panel of Fig. 18, the blue curves show the points in the Markov chain, sampling directly in $c_{1}$ vs $c_{2}$. These are the same points as those in the upper-right quadrant of the left panel of Fig. 17. In orange, we show the one-dimensional distributions of only the allowed parameter points out of all those plotted in the left panel of Fig. 18. Neither the likelihood nor the probability is parametrization invariant, so we need to apply the Jacobian
@@
Appendix B Higgs, top, di-boson and electroweak combination
In producing the global analysis of Sec. 4.4 and Fig. 14, we have combined the top sector from this paper with the previous SFitter analysis of the Higgs, di-boson, and electroweak sectors of Ref. [65], taking all data from within this reference. Note that while Ref. [65] provides constraints on Wilson coefficients in the HISZ basis; here, we provide all constraints in the Warsaw basis.
In addition to the 22 operators introduced in Sec. 2.1 and constrained by the top sector observables, a further 21 Wilson coefficients can be constrained by the addition of data from the Higgs, di-boson, and electroweak sectors. These operators are assumed to follow the same flavor symmetry conventions as introduced in Eq.(2), i.e. flavor universality applied to the first two quark generations. The notation and conventions for these 21 operators are provided in Tab. 5.
Coefficient Operator Coefficient Operator
$C_{\phi G}$ $\phi^{\dagger}\phi G_{\mu\nu}^{A}G^{A\mu\nu}$ $C_{W}$ $\varepsilon^{IJK}W_{\mu}^{I\nu}W_{\nu}^{J\rho}W_{\rho}^{K\mu}$
$C_{\phi\Box}$ $(\phi^{\dagger}\phi)\Box(\phi^{\dagger}\phi)$ $C_{d\phi,33}$ $(\phi^{\dagger}\phi)(\bar{Q}_{3}b\phi)$
$C_{u\phi,33}$ $(\phi^{\dagger}\phi)(\bar{Q}_{3}t\phi)$ $C_{\phi d}$ $\sum_{i=1}^{2}(\phi^{\dagger}i\overleftrightarrow{D_{\mu}}\phi)(\bar{d}_{i}\gamma^{\mu}d_{i})$
$C_{\phi e}$ $(\phi^{\dagger}i\overleftrightarrow{D_{\mu}}\phi)(\bar{e}\gamma^{\mu}e)$ $C_{\phi u}$ $\sum_{i=1}^{2}(\phi^{\dagger}i\overleftrightarrow{D_{\mu}}\phi)(\bar{u}_{i}\gamma^{\mu}u_{i})$
$C_{\phi b}$ $(\phi^{\dagger}i\overleftrightarrow{D}_{\mu}\phi)(\bar{b}\gamma^{\mu}b)$ $C_{\phi q}^{(3)}$ $\sum_{i=1}^{2}(\phi^{\dagger}i\overleftrightarrow{D_{\mu}}\phi)(\bar{q}_{i}\gamma^{\mu}q_{i})$
$C_{\phi D}$ $(\phi^{\dagger}D^{\mu}\phi)^{*}(\phi^{\dagger}D^{\mu}\phi)$ $C_{\phi q}^{(1)}$ $\sum_{i=1}^{2}(\phi^{\dagger}i\overleftrightarrow{D_{\mu}}\phi)(\bar{q}_{i}\tau^{I}\gamma^{\mu}q_{i})$
$C_{\phi B}$ $\phi^{\dagger}\phi B_{\mu\nu}B^{\mu\nu}$ $C_{e\phi,22}$ $(\phi^{\dagger}\phi)(\bar{l}_{2}\mu\phi)$
$C_{\phi W}$ $\phi^{\dagger}\phi W_{\mu\nu}^{I}W^{I\mu\nu}$ $C_{e\phi,33}$ $(\phi^{\dagger}\phi)(\bar{l}_{3}\tau\phi)$
$C_{\phi WB}$ $\phi^{\dagger}\tau^{I}\phi W_{\mu\nu}^{I}B^{\mu\nu}$ $C_{ll}$ $(\bar{l}\gamma_{\mu}l)(\bar{l}\gamma^{\mu}l)$
$C_{\phi l}^{(1)}$ $(\phi^{\dagger}i\overleftrightarrow{D_{\mu}}\phi)(\bar{l}\gamma^{\mu}l)$ $\text{BR}_{\text{inv}}$ invisible Higgs decays
$C_{\phi l}^{(3)}$ $(\phi^{\dagger}i\overleftrightarrow{D}^{I}_{\mu}\phi)(\bar{l}\tau^{I}\gamma^{\mu}l)$
Table 5: Additional Wilson coefficients of the Warsaw basis entering the combined analysis of the Higgs, di-boson, top, and electroweak sectors. The 21 degrees of freedom shown here are included in the global analysis of Sec. 4.4 alongside the 22 operators already constrained by top sector observables. In total, 43 coefficients are constrained in the global analysis.
Table 6 reports the numerical values of the boundaries of the 95% CL intervals shown in Fig. 14.
Coefficient Full analysis Halved theory unc. Coefficient Full analysis Halved theory unc.
$C_{\phi G}$ [-9.25, 6.35] [-5.56, 5.1] $C_{tG}$ [-0.46, 0.35] [-0.12, 0.19]
$C_{\phi\Box}$ [-1.14, 1.72] [-1.0, 1.27] $C_{Qq}^{(18)}$ [-0.29, 0.08] [-0.21, 0.08]
$C_{u\phi,33}$ [-7.03, 9.11] [-7.03, 9.11] $C_{Qq}^{(38)}$ [-0.23, 0.15] [-0.23, 0.09]
$C_{\phi e}\times 10$ [-2.73, 1.79] [-2.73, 1.59] $C_{tq}^{(8)}$ [-0.37, 0.15] [-0.26, 0.12]
$C_{\phi b}\times 10$ [-11.46, -1.25] [-10.17, -1.33] $C_{Qu}^{(8)}$ [-0.49, 0.2] [-0.43, 0.18]
$C_{\phi D}\times 10$ [-3.48, 5.4] [-3.33, 5.4] $C_{Qd}^{(8)}$ [-0.7, 0.4] [-0.68, 0.33]
$C_{\phi B}\times 10$ [-2.12, 0.63] [-2.15, 0.63] $C_{tu}^{(8)}$ [-0.38, 0.1] [-0.35, 0.1]
$C_{\phi W}\times 10$ [-3.15, 4.29] [-3.0, 4.28] $C_{td}^{(8)}$ [-0.42, 0.2] [-0.4, 0.11]
$C_{\phi WB}\times 10$ [-2.24, 1.64] [-2.24, 1.26] $C_{Qq}^{(11)}$ [-0.1, 0.06] [-0.08, 0.05]
$C_{\phi l}^{(1)}\times 10$ [-1.14, 1.07] [-1.14, 1.0] $C_{Qq}^{(31)}$ [-0.08, 0.06] [-0.06, 0.04]
$C_{\phi l}^{(3)}\times 10$ [-1.38, 0.27] [-1.36, 0.17] $C_{tq}^{(1)}$ [-0.08, 0.09] [-0.07, 0.1]
$C_{W}\times 20$ [-1.1, 1.2] [-1.1, 1.12] $C_{Qu}^{(1)}$ [-0.08, 0.11] [-0.09, 0.1]
$C_{d\phi,33}\times 20$ [-0.94, 1.51] [-0.58, 1.36] $C_{Qd}^{(1)}$ [-0.15, 0.14] [-0.13, 0.1]
$C_{\phi d}\times 20$ [-2.83, 3.81] [-2.19, 3.05] $C_{tu}^{(1)}$ [-0.11, 0.08] [-0.1, 0.09]
$C_{\phi u}\times 20$ [-1.75, 1.39] [-1.75, 1.31] $C_{td}^{(1)}$ [-0.14, 0.12] [-0.14, 0.11]
$C_{\phi q}^{(3)}\times 20$ [-1.56, 0.84] [-1.54, 0.8] $C_{\phi Q}^{(3)}$ [-0.66, 0.32] [-0.56, 0.23]
$C_{\phi q}^{(1)}\times 20$ [-1.39, 1.08] [-1.46, 1.08] $C_{tW}$ [-0.16, 0.31] [-0.13, 0.26]
$C_{e\phi,22}\times 100$ [-0.29, 0.58] [-0.3, 0.58] $C_{bW}/10$ [-0.17, 0.19] [-0.12, 0.12]
$C_{e\phi,33}\times 100$ [-1.35, 2.06] [-1.26, 1.37] $C_{tZ}/10$ [-0.3, 0.17] [-0.24, 0.09]
$C_{ll}\times 100$ [-4.61, 0.21] [-4.51, 0.0] $C_{\phi Q}^{(1)}/10$ [-0.56, 0.78] [-0.41, 0.79]
$\text{BR}_{\text{inv}}$ [0, 7.6] [0, 7.03] $C_{\phi tb}/10$ [-0.54, 0.48] [-0.37, 0.36]
$C_{\phi t}/100$ [-0.2, 0.11] [-0.17, 0.07]
Table 6: Numerical values for the 95% CL limits shown in Fig. 14. We emphasize that the reduction of the theory uncertainties by a factor two is entirely hypothetical.